# RealitySummary: Exploring On-Demand Mixed Reality Text Summarization and Question Answering using Large Language Models

### v2

@@ 6.1. Implementation
The final version of RealitySummary builds on the earlier Vision Pro implementation with two key improvements: 1) Users can now add multiple draggable cards on-demand by either making a new query or clicking an “add” button. User can spatially anchor these cards anywhere in their immediate environment and can also re-select cards to update them with new prompts or text from the scene. These cards are stored as JSON objects in a real-time database. 2) The system allows users to mirror their device screen within the Vision Pro for high-resolution viewing. This functionality — implemented through a WebRTC video call or via Mac Virtual Display on the Vision Pro — lets users view their external devices at full resolution in the MR environment and also allows RealitySummary to more accurately capture and process on-screen content.
6.2. Method
Five authors (A1-A5) used the system over the course of 3 weeks, integrating it into their day-to-day activities and accumulating 30 total hours of use. Due to physical fatigue and battery limitations of the Vision Pro, each session lasted less than 2 hours. The authors independently documented their use cases, insights, frustrations, thoughts, and reflections while using RealitySummary, which we later thematically coded. During the study, each participating author recorded their observations via a phone microphone, following a think-aloud process. Additionally, the first author conducted follow-up interviews to gain deeper insights into their experiences.
@@ 6.3. Results
Throughout the diary study, participants used the system in various settings, including labs, kitchens, outdoor areas, and offices, though the majority of usage took place at desks in work environments. They engaged with a wide range of reading materials, such as academic papers, news articles, code, emails, calendars, Google Scholar, classroom assignments, lecture slides, textbooks, and YouTube videos. Participants quickly adapted the system to fit a variety of personal, academic, and technical tasks. For example, A4 used the assistant not only to summarize email content but also to gather higher-level insights, such as identifying how many emails were sent by a particular sender. A2 utilized the system to support programming assignments, where they used the system to break down complex coding problems into manageable steps and provided detailed explanations of specific code snippets. A2 also leveraged the system for learning aids, such as generating to-do lists and quiz questions based on problem descriptions. The ability to spatially organize tasks and concepts allowed him to effectively manage his workload, maintain focus, and dynamically track his progress. Additionally, in his language learning efforts, A2 used the assistant to generate contextual vocabulary lists from Japanese reading materials, which helped him focus on the most relevant and frequently used words. A5 employed it to extract a step-by-step sequence of keyboard commands from an article on rebooting a MacBook. Overall, the diary study highlights the system’s flexibility in supporting diverse tasks, from email management and coding to technical troubleshooting, language learning, and academic research.
6.4. Insights
@@
From these experiences, participants shared various insights, frustrations, and potential future improvements for the current prototype. Our diary study also provided insight into the use of an always-on reading assistant in more realistic work-oriented settings.
Quick and Simple Summary Responses Meet Many Immediate Needs
@@
Participants valued the system’s ability to deliver quick, concise responses, which were often sufficient for immediate clarifications. A2, for instance, appreciated how the system provided brief answers when they needed a rapid comparison of philosophical concepts, like distinguishing between act and rule utilitarianism. In many cases, just a high-level summary captured through manual input was enough to satisfy their needs. A4 found this method to be a useful shortcut, as pressing a button was faster and less disruptive than speaking and often provided exactly the information they were seeking. These simple, proactive summaries were particularly beneficial for tasks that required quick comprehension and minimal interruption.
Minimal Context Switching Enhances Focus and Flow
Participants appreciated the integration of LLM into the MR interfaces, in part, because it helped them minimize context switching, allowing them to maintain greater focus. The system’s automatic text extraction feature enabled participants to stay in a flow state. For example, A2 noted that the interface allowed them to concentrate on their work without the mental disruption of switching between tasks or manually inputting context, even when using a computer screen. This seamless workflow kept participants engaged, enhancing productivity and reducing cognitive load. A1 highlighted how the system functioned like a “teacher” by offering real-time assistance without requiring them to step away from their tasks. Participants also remarked that on-demand support boosted their efficiency during complex tasks like coding or writing, where focus is critical.
@@
Participants highlighted the advantages of peripheral output modalities. For instance, A2 used the system as a reading aid while learning Japanese. Instead of frequently opening a dictionary, which disrupted the reading flow, they generated vocabulary flashcards around the document. This provided the needed language support without interrupting the reading and learning process. A4 also noted that they would often cluster and group these generated cards thematically to the sides of the display for quick reference and then delete them once they were no longer needed. Similarly, A1 used the system to “pin” generated code versions and key concepts around their workspace, allowing for easy access without breaking focus and enabling them to quickly revisit important points as needed.
Spatial Organization Improves Recall and Multitasking
A key finding from the study was how participants leveraged the system’s spatial organization features, similar to arranging post-it notes on a board. Both A1 and A2 utilized this capability to group related concepts together. For example, A2 clustered cards on CPU architecture, placing related topics like buses and registers nearby, much like organizing post-its by theme. This spatial arrangement enabled quick access and improved recall, as users could visually track where specific information was placed. Similarly, A1 pinned cards documenting different problem-solving methods across their workspace, ensuring easy access when needed. This post-it-like organization supported memory retention and task management by allowing users to arrange information in a way that aligned with their thought processes.
6.5. Limitations
Challenges Controlling What to Capture
Currently, the system captures the user’s entire field of view, but participants expressed a desire for more control over what is captured. For example, A1 struggled to get the system to ignore personal emails displayed next to the article he was reading, leading to summaries that inadvertently included private information. Similarly, A2 wanted to focus on a specific code snippet to obtain more accurate results while debugging. Additionally, the lack of visual feedback left users uncertain about what OCR content or context was being processed in each request. To address these issues, the system could offer enhanced visual feedback and allow users to control specific areas of the screen, such as through gaze interaction, which would help mitigate privacy concerns and prevent the inclusion of irrelevant or personal content.
Cumbersome Re-Prompting Process for Common and Repeated Tasks
Participants frequently found themselves re-prompting similar queries for routine tasks, such as requesting bullet point summaries or filtering by specific criteria. A potential improvement could be the addition of dynamic buttons that save commonly used prompt types, allowing users to easily reuse them later. Additionally, participants expressed frustration with testing out lots of different prompts to generate a complex output, such as aggregating the Google Scholar page into a timeline. Saving or sharing the same specific instructions for repeated tasks might mitigate these problems.
Managing Growing Numbers of Summaries
In the first prototype, participants found it distracting to have multiple summary cards anchored to the image in their workspace. However, they did not find this frustrating when they had control over the placement and could easily delete the cards. Additionally, participants noted that the system could sometimes generate an overwhelming number of cards. For instance, A2 mentioned that while analyzing a poem, the system created a card for each sentence and its meaning, making it difficult to manage information overload. This feedback suggests that there is potential value in exploring new interactive and semi-automated methods for efficiently sorting, snapping, organizing, filtering, and removing these cards. Such improvements could enhance the usability of the system and help users better manage the content generated.
@@ 7.1. Unique Benefits of Combining MR and AI
The always-on camera in MR enables implicit input, distinguishing it from the explicit input required in desktop interfaces, such as opening an app, typing queries, or copy-pasting contextual text. Participants in in-the-wild and diary studies appreciated the system’s ability to capture information without requiring explicit effort. This always-on input allows for hands-free exploration in both spatial and temporal contexts, such as scanning entire bookshelves as a query or utilizing the history of flipped pages. Implicit input reduces cognitive load and minimizes context switching, as users are not required to actively specify what information should be used for the LLM’s context.
@@
We found that MR’s spatial output allows information to be presented in a peripheral, background manner. In contrast, other output modalities, such as audio or screen-based displays, often require the user’s foreground attention to consume generated content, disrupting focus and interrupting workflow. With MR and peripheral outputs, however, AI-generated content becomes less intrusive, as it can be easily ignored if not needed, simply by choosing not to look at it. For example, participants in the diary study often appreciated this feature for learning aids, such as understanding textbooks or accessing always-on vocabulary references when learning a foreign language. This peripheral output is a unique aspect of the integration of MR and AI. We believe this capability would promote more calm human-AI interactions in the future, reducing disruptions to focused attention by combining implicit input with unobtrusive background support.
7.2. Emergent Use Cases for Mixed Reality Text Augmentation
Both the in-the-wild and diary studies revealed a variety of interesting emergent use cases. To provide a more comprehensive view of how users engage with these systems, we have categorized these into seven groups:
@@ 7.2. Emergent Use Cases for Mixed Reality Text Augmentation
Participants also utilized the system to extract key information, which can be translated into a more consumable format. This aligns with design elicitation results, which suggest tables of contents, keyword lists, timelines, and comparison tables. For example, one participant extracted vocabulary flashcards to assist with foreign language reading, while another participant created a table of keyboard shortcut commands from a troubleshooting guide to use as a reference. In these examples, the consumable format, such as tables or vocabulary lists, serves as helpful references for their tasks.
@@
Participants frequently used extracted text as a query to retrieve additional information from external sources. Echoing the information types identified in our elicitation study, participants queried specific names or keywords from their reading materials. For instance, one participant retrieved details about a player, team, and standings from a sports-related article, while another fact-checked claims on-demand when reading news articles.
@@
Beyond the above use cases identified in the elicitation study, participants frequently asked for recommendations or suggestions based on the current context. For instance, users asked for book recommendations when looking at a bookshelf. Beyond the reading context, one participant asked for advice on which trash bin can be used based on disposal instructions, while others inquired about healthier menu options or products. They also asked for recommendations from external sources. For instance, participants asked for suggestions for relevant research papers when reading academic articles.
@@
An interesting use case involved using the system to aggregate high-level information. Unlike simple information extraction, aggregation often involves operations like filtering and counting. For example, some participants used the system to count emails based on specific filters, while others filtered and counted first-author papers from a researcher’s Google Scholar page. This allowed users to perform quick, approximate queries that would normally require more complex searches.
@@
Participants also used the system to transform their reading content into interactive, actionable outputs. For instance, some participants applied our system to generate quizzes from class notes or textbooks, while others decomposed high-level assignments into step-by-step to-do lists. One participant asked the system to create reminder cards that remained spatially persistent. Participants also use our system to translate from one language to another. Unlike information extraction, these tasks involved converting content into actionable, interactive outputs.
@@
These emergent use cases highlight novel ways of interacting with AI-generated information, allowing users to engage with individual snippets and convert them into dynamic outputs. They also illustrate how participants adapted the assistant to handle different types of content, which suggests a broad space of possible designs for such systems.
7.3. Future Opportunities and Design Implications
As both MR and LLM technologies continue to evolve rapidly, considerable opportunities exist for tools that more richly integrate them. Based on our three studies, we identify several key opportunities for future research and design.
Enabling Greater Control Over Capture
While the system’s implicit input offers significant benefits, it also raises privacy concerns due to the automatic capturing without the user’s constant awareness. To address this, future designs should give users control over what data is captured. As noted in the diary study, participants expressed a desire for more transparency and control over the captured content. One approach could involve providing visual feedback, allowing users to see and select the specific areas or information being captured. Future systems should balance the benefits of implicit input with privacy considerations by offering user control and visual feedback.
Improving Organization and Filtering of Interactive Cards
Participants expressed the need for more efficient ways to manage the growing number of spatially organized content cards. While some diary study participants manually arranged these cards around documents, others suggested more advanced organizational tools. Features like sorting, snapping, filtering, or automatically organizing cards could simplify workflows and reduce clutter. Such capabilities would be particularly beneficial for long-term tasks, where the accumulation of content can become overwhelming. Integrating these organizational tools would enhance usability, supporting users in managing dynamic and task-specific information over time, which enables better workspace management for daily tasks.
Balancing Between Reusable Prompts and On-Demand Custom Queries
Our studies highlighted the importance of accommodating both reusable prompts and custom queries. Each has its advantages: prepared prompts can generate complex outputs such as timelines or comparison tables, while custom queries offer users flexibility in generating content. The diary study participants expressed a desire for dynamic interfaces that could quickly reuse and adapt prompts to different contexts. For example, some suggested having customizable buttons that automatically generate summaries or lists in a consistent style. Others mentioned the need for a more efficient way to reuse queries across tasks, such as repeatedly generating lists or comparisons for class topics. This feedback highlights the importance of designing adaptive interfaces that allow users to seamlessly reuse and modify their prompts, improving efficiency and user satisfaction.
Facilitating Seamless Transitions Between Screen and Space
Participants also valued the ability to drag content cards from the system into their computer, such as pasting generated answers directly into coding tasks. Some participants also envisioned this feature being useful for writing tasks, allowing for a smooth transfer of generated content. Future systems should explore leveraging cross-device interactions between MR and other devices to facilitate seamless transitions, enhancing the workflow between screen-based and spatial interactions.
Expanding Beyond Text Input to Capture Images and Everyday Objects
Participants expressed a strong desire for the system to recognize not only text but also images. For example, the in-the-wild study revealed the limitations of inquiring about a diagram that the system could not recognize. Users also anticipated broader applications beyond text recognition, such as generating recipes based on ingredients in the fridge or providing information on real-world objects. Enhancing the system with multi-modal LLM capabilities could address these needs, expanding its utility beyond text-based interactions.
Personalization Through Always-On History
Lastly, participants showed interest in an “always-on” history feature that would allow them to recall previous interactions with the system, creating a more personalized experience. They speculated that such a system could eventually serve as a personalized assistant, continuously capturing and understanding context through an always-on camera. Although current hardware limitations, such as battery life and physical fatigue, prevent continuous long-term use, we believe future MR interfaces will seamlessly integrate with AI to provide both context-aware and personalized assistance based on users’ daily activity histories.

### v3

@@ 4.4. RS-Wild Observations
4.5. Opportunities for the Third Iteration
Overall, the design of the second prototype enhanced its capacity for real-world use. Compared to the earlier HoloLens 2 prototype, the Vision Pro’s superior rendering quality and broader field of view provided a better reader experience. Additionally, the simple system design greatly improved usability despite the absence of document tracking. While RS-Wild showed great potential, we also identified several feature-level opportunities for improvement.
Sustained History and Recall
Participants reported that the minimalist design effectively reduced visual clutter. However, it also limited the system’s overall utility. Although readers appreciated the simple interface, many envisioned options to save or revisit previously-generated responses. Several participants suggested that a history feature or the ability to store cards for future reference would substantially enhance the reader experience.
Enhanced Screen Viewing for Extended Use
Participants highlighted the potential for improving the experience of viewing phone and computer screens through the Vision Pro, particularly for longer viewing periods. While the Vision Pro’s video passthrough worked well for short-term tasks like quickly reviewing an assignment, participants envisioned improvements that could better support extended work sessions, especially when interacting with computing devices. Mirroring the reader’s screen(s) visually in high resolution through the headset would further align the system with the demands of realistic, work-related tasks in everyday contexts.
Our in-the-wild study was also inherently limited by its short-term, public use — with each session lasting only one hour per participant. As a result, participants did not have the opportunity to fully engage with the system in their personal workflows. For instance, due to privacy concerns related to screen recording, participants rarely used our system for more personalized content, such as emails, messages, or calendars. Furthermore, since we encouraged participants to explore multiple scenarios, they tended to experiment broadly rather than focusing on any specific task in depth. While the in-the-wild study provided valuable insights into real-world use cases, it only captured short-term interactions in public spaces. There is still much to learn about the sustained, regular use of reading assistants in daily life, particularly in how readers interact with information over time, revisit previous queries, or support ongoing tasks.
5. RS-Diary: Diary Study with the Third Prototype
To explore more realistic longer-term use cases, we developed a third prototype variant (RS-Diary) incorporating feedback from the in-the-wild study. This version (Figure 4) introduced the ability to create, retain, and organize multiple information cards, along with support for mirroring content from external devices directly to the Vision Pro.
@@ 5.1. Implementation
RS-Diary builds on the earlier Vision Pro implementation with two key improvements: 1) Readers can now add multiple draggable cards on-demand by either making a new query or clicking an “add” button (which spawns a new summary card based on the document content currently in front of them). They can also spatially anchor these cards anywhere in their immediate environment and can re-select existing cards to update them with new prompts or text content. These cards are stored as JSON objects in a real-time database. 2) The system also allows readers to mirror their device screen to the Vision Pro for high-resolution viewing. This functionality—implemented through a WebRTC video call or via Mac Virtual Display on the Vision Pro—lets readers view their computer or tablet screens at full resolution in the MR environment and also allows RS-Diary to more accurately capture and process on-screen content. This increased fidelity allows readers to perform a wide variety of everyday reading, coding, browsing, and watching activities for extended periods without the legibility and eye strain experienced with video passthrough.
5.2. Study Method
Four authors and collaborators (AA, AM, AJ, AW, and CT  [^2] The first, fourth, fifth, and sixth authors of this submission and one external acknowledged collaborator. The second letter of each code corresponds to the author/collaborator’s first initial.) used the system over the course of 3 weeks, integrating it into their day-to-day activities and accumulating 30 total hours of use. Due to physical fatigue and battery limitations of the Vision Pro, each session lasted less than 2 hours. The participants independently documented their use cases, insights, frustrations, thoughts, and reflections while using RS-Diary. During the study, each participant recorded their observations via a smartphone audio recorder and a personal reflections diary for notes, following a think-aloud process. Additionally, the first author conducted reflective debrief conversations with the other four participants at the end of the study to gain deeper insights into their experiences. We then conducted a reflexive thematic analysis (Clarke and Braun, 2014) using transcripts of these discussions, from which we identified and refined key observations.
@@ 5.3. Results
Throughout the diary study, participants used the system in various settings, including labs, kitchens, outdoor areas, and offices, though the majority of usage took place at desks in office environments. They engaged with a wide range of reading materials, such as academic papers, news articles, code, emails, calendars, Google Scholar, classroom assignments, lecture slides, textbooks, and YouTube videos. Participants quickly adapted the system to fit a variety of personal, academic, and technical tasks. For example, AW used the assistant not only to summarize email content but also to gather higher-level insights, such as identifying how many emails were sent by a particular sender. AM utilized the system to support programming assignments, where they used the system to break down complex coding problems into manageable steps and provided detailed explanations of specific code snippets. AM also leveraged the system to create learning aids, such as generating to-do lists and quiz questions based on problem descriptions. The ability to spatially organize tasks and concepts allowed him to balance his workload, maintain focus, and dynamically track his progress. Additionally, in his language learning efforts, AM used the assistant to generate contextual vocabulary lists from Japanese reading materials, which helped him focus on the most relevant and frequently used words. CT employed it to extract a step-by-step sequence of keyboard commands from an article on rebooting a MacBook. Overall, the diary study highlighted the system’s flexibility in supporting diverse tasks, from email management and coding to technical troubleshooting, language learning, and academic research.
5.4. RS-Diary Observations
@@
From these experiences, participants shared various thoughts, frustrations, and potential future improvements. Our diary study also provided insight into the use of an always-on reading assistant in more realistic work-oriented settings.
Quick Summary Responses Meet Many Immediate Needs
@@
Participants valued the system’s ability to deliver quick, concise responses, which were often sufficient for immediate clarifications. AM, for instance, appreciated how the system provided brief answers when they needed a rapid comparison of philosophical concepts, like distinguishing between act and rule utilitarianism. In many cases, just a high-level summary of the page in view was enough to satisfy their needs. AW found this method to be a useful shortcut, as pressing the virtual “add” button was faster and less disruptive than speaking and often provided exactly the information they were seeking. These simple, proactive summaries were particularly beneficial for tasks that required quick comprehension and minimal interruption.
@@
Participants highlighted the advantages of cards that could exist in the periphery of their current workspace. For instance, AM used the system as a reading aid while learning Japanese. Instead of frequently opening a dictionary, which disrupted the reading flow, they generated vocabulary flashcards around the document. This provided language support without interrupting their reading and learning process. AW also noted that they would often cluster and group these generated cards thematically to the sides of the display for quick reference and then delete them once they were no longer needed. Similarly, AA used the system to “pin” generated code versions and key concepts around their workspace, allowing for easy access without breaking focus and helping them quickly revisit important points as needed.
Spatial Organization for Recall and Multitasking
Perhaps the biggest emergent pattern in the study was how all participants leveraged the system’s spatial organization features, similar to arranging post-it notes on a board. Both AA and AM utilized this capability to group related results and summaries together. For example, AM clustered cards on CPU architecture, placing related topics like buses and registers nearby, much like organizing post-its by theme. This spatial arrangement enabled quick access and improved recall, as readers could visually track where they had placed specific information. Similarly, AA pinned cards documenting different problem-solving methods across their workspace, ensuring easy access when needed. These organizational structures supported memory retention and task management by allowing readers to arrange information in a way that aligned with their thought processes.
@@ 6. Lessons Learned from Three Studies
6.1. Key Insights from the Three Studies
Minimal Context Switching Enhances Focus and Flow
Participants across all three studies appreciated the integration of LLM into the MR interfaces, in part, because it helped them minimize context switching, allowing them to maintain greater focus. Several participants reported that the system’s automatic text extraction feature helped them stay in a flow state. For example, AM noted that the interface allowed them to concentrate on their work without the mental disruption of switching between tasks or manually inputting context, even when using a computer screen. In particular, members of the author team in the RS-Diary study remarked on how the seamless workflow kept them engaged, enhancing productivity and reducing cognitive load. AA highlighted how the system functioned like a “teacher” by offering real-time assistance without requiring them to step away from their tasks. Participants also remarked that on-demand support boosted their efficiency during complex tasks like coding or writing, where focus is critical.
A key benefit of this workflow is its support for implicit inputs, which dramatically lower the barrier to asking questions of LLMs. This automatic, continuous content capture broadens the range of applications beyond what participants initially envisioned. For instance, W4 mentioned they had never thought to type a question into ChatGPT when looking at a product label. Even for common reading tasks, they appreciated that the summary was automatically generated and updated in the background as they flipped through pages in a library. Overall, the always-on, implicit assistant made it effortless to switch context and maintain focus.
Potential Risks of AI-Driven Augmented Reading
Participants also expressed several concerns about potential risks, notably privacy issues due to the possibility of constant surveillance. Despite this concern, most participants expressed interest in continue to use these kinds of systems, provided they could control when it was active. For example, W1 expressed the desire to capture all text throughout her entire life, allowing for an extensive searchable archive of everything she has encountered. Such use cases, however, also pose social privacy questions—particularly when systems might record other individuals or their content.
Additionally, participants mentioned the risk of decreased motivation and skill development. For example, W6 mentioned that relying too heavily on the system for translations could reduce the motivation to learn new languages. Similarly, W11 worried that immediate access to answers for every question might contribute to a broader intellectual complacency. Overall, participants wanted the flexibility to choose when and how the system would be used to mitigate these risks.
Trust of AI-Generated Content
Recognizing the known shortcomings or current LLMs, participants also approached the system’s responses with caution, often verifying the answers by reading the source materials and highlighting the necessity of consulting additional sources. There was a general agreement among participants in all studies that they would not fully trust the system’s responses to be complete or correct. Conversely, we noted some intriguing discussions regarding how participants perceived our system, either as a mere tool or more like a personal companion. While most participants engaged with the system as a conventional chatbot, some regarded it more as a personal companion than a mere tool. Notably, W3 and W5 described the system as a friend for discussing and inquiring about their readings. The system’s ability to provide always-on, immediate responses facilitates this personification.
@@ 6.2. Unique Benefits of Combining MR and AI
The always-on camera in MR enables implicit input, distinguishing it from the explicit input required in desktop interfaces, such as opening an app, typing queries, or copy-pasting contextual text. Participants in in-the-wild and diary studies appreciated the system’s ability to capture information without requiring explicit effort. This always-on input allows for hands-free exploration in both spatial and temporal contexts, such as scanning entire bookshelves as a query or utilizing the history of flipped pages. Implicit input reduces cognitive load and minimizes context switching, as readers are not required to actively specify what information should be used for the LLM’s context.
@@
We found that MR’s spatial output allowed information to be presented in a peripheral, background manner. In contrast, other output modalities, such as audio or screen-based visuals, often require the reader’s foreground attention to consume generated content, disrupting focus and interrupting workflows. With MR and peripheral outputs, however, AI-generated content becomes less intrusive, as it can be easily ignored if not needed, simply by choosing not to look at it. For example, participants in the diary study often appreciated this feature for learning aids, such as understanding textbooks or accessing always-on vocabulary references when learning a foreign language. This peripheral output is a unique aspect of the integration of MR and AI. We believe this capability would promote more calm human-AI interactions in the future, reducing disruptions to focused attention by combining implicit input with unobtrusive background support.
6.3. Emergent Use Cases
Both the in-the-wild and diary studies revealed a variety of interesting emergent use cases for MR text augmentation. To give a more comprehensive view of how readers might use these systems, we identify seven distinct approaches:
@@ 6.3. Emergent Use Cases
Participants also utilized the system to extract key information, which can be translated into more consumable formats like tables of contents, keyword lists, timelines, and comparison tables. For example, one participant extracted vocabulary flashcards to assist with foreign language reading, while another participant created a table of keyboard shortcut commands from a troubleshooting guide to use as a reference. In these examples, the consumable format, such as a table or vocabulary list, serves as helpful references for their tasks.
@@
Participants frequently used extracted text as a query to retrieve additional information from external sources. Participants often queried specific names or keywords from their reading materials. For instance, one participant retrieved details about players, teams, and standings from a sports-related article, while another fact-checked claims on-demand when reading news articles.
@@
Participants frequently asked for recommendations or suggestions based on the current context—for instance, asking for book recommendations when looking at a bookshelf. Outside the context of reading, one participant asked for advice on which trash bin to use based on the disposal instructions on an item, while others inquired about healthier menu options or products. They also asked for recommendations from external sources, including asking for related research papers when reading academic articles.
@@
An interesting use case involved using the system to aggregate high-level information. Unlike simple information extraction, aggregation often involves operations like filtering and counting. For example, some participants used the system to count emails based on specific filters, while others filtered and counted first-author papers from a researcher’s Google Scholar page. This allowed readers to perform quick, approximate queries that would normally require more complex searches.
@@
Participants also used the system to transform their reading content into interactive, actionable outputs. For instance, some participants applied our system to generate quizzes from class notes or textbooks, while others decomposed high-level assignments into step-by-step to-do lists. One participant asked the system to create reminder cards that remained spatially persistent. Participants also used our system to translate from one language to another. Unlike information extraction, these tasks involved converting content into actionable, interactive outputs.